Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS June 2009, Page 2 of 6  Not logged in ELOG logo
ID Date Author Project Subject Run Number Events StartTime EndTime Text Attachments
  88   Mon Aug 24 15:28:49 2009 Georg AuzingerPt_ITE512Pt_ITE512 7deg airrun02820k24.08.09 15:39:5024.08.09 16:40:54Pt: 7 Grad, HV=200V, 0% RH, beam on, 120GeV
ITE512: HV=70V,
X-Y-Table:
x=99.6, y=36
  
  87   Mon Aug 24 14:23:01 2009 Georg AuzingerPt_ITE512Pt_ITE512 5deg airrun02720k24.08.2009 14:27:0724.08.2009 15:26:20Pt: 5 Grad, HV=200V, 0% RH, beam on, 120GeV
ITE512: HV=70V,
X-Y-Table:
x=99.6, y=36
  
  86   Mon Aug 24 13:01:14 2009 Georg AuzingerPt_ITE512Pt_ITE512 3deg airrun02620k24.08.09 11:24:3324.08.2009 14:11:45Pt: 3 Grad, HV=200V, 0% RH, beam on, 120GeV
ITE512: HV=70V,
X-Y-Table:
x=99.6, y=36
  
  85   Mon Aug 24 11:58:23 2009 Georg AuzingerPt_ITE512Pt_ITE512 1deg airrun02520k24.08.09 11:24:3324.08.09 12:40:25Pt: 1 Grad, HV=200V, 0% RH, beam on, 120GeV
ITE512: HV=70V,
X-Y-Table:
x=99.6, y=36
  
  84   Sun Aug 23 23:38:27 2009 Georg Auzinger Phone numbers     Georg #70
Wolfgang - Building: 41, Room: 204,
Tel: ?
  
  83   Sun Aug 23 17:19:50 2009 Christian IrmlerPt_ITE512internal calibrationcal011 23.08.2009 16:30:4323.08.2009 17:02:14Pt: 1 Grad, HV=200V, 0% RH, intcal, beam
off
ITE512: HV=70V,
  
  82   Sun Aug 23 17:18:25 2009 Christian IrmlerPt_ITE512pedestal correctionpedcorr008 23.08.2009 17:17:36  Pt: 1 Grad, HV=200V, 0% RH, intcal, beam
off
ITE512: HV=70V,
  
  81   Sun Aug 23 14:03:14 2009 Manfred ValentanPt_ITE512Pt & ITE512 center position, 1 degrun0242000023.08.2009 13:59:06 Pt: 1 Grad, HV=200V, 0% RH
ITE512: HV=70V
X-Y-Table: x=99.6, y=36
  
  80   Sun Aug 23 13:15:42 2009 Marko DragicevicPt_ITE512Current and Humindity during first Run    Current of Pt modules during first run
Data
of the run was dicarded.
Compressed air flow was at ~2 bars
 IandRH_first_Pt_run.pngPicture_1.png 
  79   Sun Aug 23 08:13:31 2009 Thomas BergauerITE1282nd run with ITE128: top position (4mm below center)run023100k23.08.2009 08:06:5823.08.2009 09:30:53ITE128
HV = 70V, 40MHz, 120 GeV
TLU trigger w telescope 
  
  78   Sun Aug 23 02:18:20 2009 Thomas BergauerITE1282nd run with ITE128: top position (3mm below center); during nightrun02240789523.08.2009 02:16:2123.08.2009 08:04:09ITE128
HV = 70V, 40MHz, 120 GeV
TLU trigger w telescope 
 Picture_1.png 
  77   Sun Aug 23 01:46:03 2009 Thomas BergauerITE1282nd run with ITE128: central positionrun021100k23.08.09 00:10 (ca.)23.08.2009 02:15:10ITE128
HV = 70V, 40MHz, 120 GeV
TLU trigger w telescope 
 Picture_1.png 
  76   Sat Aug 22 23:06:22 2009 Christian IrmlerBELLEpedestal correction 40 MHz transparent mode + coolingpedcorr007 22.08.2009 23:04:2922.08.2009 23:13:51Belle DUT, HV=80V, 40 MHz, beam off
1 sample, transparent mode
pedestal correction
  
  75   Sat Aug 22 22:14:50 2009 Christian IrmlerBELLEinternal calibration 40 MHz transparent mode + coolingcal010 22.08.2009 22:24:22 Belle DUT, intcal
HV=80V, 40 MHz, beam off
1 sample, transparent mode
  
  74   Sat Aug 22 19:19:20 2009 Christian IrmlerBELLEAPVDAQ + telescope 40 MHz transparent mode + coolingrun02075k22.08.2009 19:17:2922.08.2009 22:09:53Belle DUT, HV=80V, 40 MHz, 120 GeV
6 samples, transparent mode
chiller set to 13 deg,  2900
  
  73   Sat Aug 22 15:52:33 2009 Christian IrmlerBELLEAPVDAQ + telescope 31.8 MHz transparent mode + coolingrun01975k22.08.2009 15:50:4022.08.2009 19:14:21Belle DUT, HV=80V, 31.8 MHz, 120 GeV
6 samples, transparent mode
chiller set to 13 deg,  2900
  
  72   Sat Aug 22 15:28:00 2009 Christian IrmlerBELLEpedestal correction 31.8 MHz transparent mode + coolingpedcorr006 22.08.2009 15:25:3522.08.2009 15:36:00Belle DUT, HV=80V, 31.8 MHz, beam off
1 sample, transparent mode
pedestal correction
  
  71   Sat Aug 22 15:12:44 2009 Thomas Bergauercommonbeam profile    playing with TRIM4 in vertical
plane (setting to 20A current) shifted beamspot
back to center position
 20090822_1511_beam_conditions.jpg 
  70   Sat Aug 22 14:46:12 2009 Christian IrmlerBELLEinternal calibration 31.8 MHz transparent mode + coolingcal009 22.08.2009 14:43:5522.08.2009 15:22:37Belle DUT, intcal
HV=80V, 31.8 MHz, beam off
1 sample, transparent mode
  
  69   Sat Aug 22 13:34:14 2009 Christian IrmlerBELLEpedestal correction 31.8 MHz transparent mode (cooling?)pedcorr005 22.08.2009 13:31:2722.08.2009 13:41:42Belle DUT, intcal
HV=80V, 31.8 MHz, beam off
1 sample, transparent mode
  
ELOG V3.1.4-966e3dd